### v1

@@ 3.2 Color-Magnitude Diagram (CMD) Matching with KDTree
3.3 Principal Component Analysis (PCA) and DBSCAN Clustering
@@
Figure 2: Spatial distribution of the identified core members (blue dots) for the five clusters with their detected tidal tails (red and coral dots), the N-body simulated cluster (empty dots), the Jacobi limit (dotted black circle), and the expected orbit (gray arrow) pointing towards the direction of motion (Refer to Subsection 4.2). Crosses indicate manually removed stars for NGC 2281.
@@ 3.4 Tangential Projection of the Celestial Coordinates
We employed a tangential projection of celestial coordinates onto a 2D Cartesian plane to analyze the spatial distribution of tidal tails in star clusters. This approach simplified visualization and enabled accurate spatial analyses by mitigating the effects of spherical geometry. The celestial coordinates of stars are expressed in right ascension ($\alpha$) and declination ($\delta$), both in degrees. To project these spherical coordinates onto a tangent plane, we define a reference point, $(\alpha_{0},\delta_{0})$, also in degrees, corresponding to the mean position of the cluster core. Using this reference, the transformation to the tangent plane is given by:
@@
Although the projection provides a convenient 2D representation, it inevitably reduces information from the full 3D structure.
@@ 3.5 Beyond the Jacobi Radius
The final criterion for checking our detection of the tidal tail was the spatial location of the stars relative to the best-fit Jacobi radius values for each cluster taken from Hunt & Reffert (2024) as provided in Table 1. Beyond this radius, stars are more likely to be stripped from the cluster and form tidal tails. The Jacobi radius ($r_{J}$, in pc) is calculated using the following formula:
$r_{J}=R_{\text{gc}}\left(\frac{M_{c}}{2M_{g}}\right)^{1/3}$ (5)
where $R_{\text{gc}}$ is the galactocentric distance of the cluster (in pc), $M_{c}$ is the mass of the cluster (in $M_{\odot}$), and $M_{g}$ is the galactic mass enclosed within $R_{\text{gc}}$ (also in $M_{\odot}$). This formula gives us an estimate of the region beyond which stars are more likely to be influenced by the galactic tidal field. We note that $r_{J}$ is not used as a strict boundary in Figure 2 for defining tidal tail stars. Due to projection effects in the sky plane, some tail stars may still appear within the projected $r_{J}$. Our use of this limit is only to show that the detected features extend beyond the theoretical boundary.
Table 2: Columns 2–4: Galactocentric positions of the identified members of the 5 open clusters; Column 5: Best fitting Jacobi radius from Hunt & Reffert (2024); Columns 6–7: Approximate projected extent of the tidal tails beyond the Jacobi Radius determined using the angular separation from the cluster center and the cluster distance, with the orbit direction aiding in identifying the tail components (Subsection 4.2); Columns 8-9: Number of stars in leading and trailing tails; Columns 10-12: Binary fractions (for mass ratio >0.5>0.5) in the cluster, leading tail and trailing tail (Subsection 4.3).
Name X Y Z $r_{J}$ $L_{lead}$ $L_{trail}$ $N_{lead}$ $N_{trail}$ $BF_{cluster}$ $BF_{lead}$ $BF_{trail}$
@@ 3.6 Luminosity Function
The luminosity functions (LFs) for each cluster and its tidal tails were constructed to facilitate comparison, which can be seen in the subplot (h) of plots 7 to 11 in the appendix. Specifically, LFs were generated for two distinct regions: the identified core cluster, and the detected tidal tails, including the leading tail and the trailing tail, each using a consistent bin size $M_{G}$ of 1 mag. The faintest one or two bins in both regions for every cluster contained fewer counts, probably due to observational incompleteness and stellar evaporation. These bins were excluded from the linear fit, which was performed using the following relation:
$\log_{10}N\propto a\times M_{G}$ (6)
@@
4 Tidal Tail Analysis
We analyzed the kinematic properties of each cluster and its tidal tails using proper motion vectors and orbit integration. Proper motions were used to characterize relative stellar motions and identify subtle patterns, while orbit integration traced past and future trajectories within the Milky Way’s potential. We also determined and analyzed the luminosity functions and binary fractions for both clusters and their tails.
4.1 Correction of PM Vectors for Projection Effects
Projection effects influence the apparent proper motions and radial velocities of individual stars within a star cluster due to the relative position and motion of the cluster with respect to the Sun (van Leeuwen, 2009; Jadhav et al., 2024). These effects arise because of the angular separation of stars from the cluster center and the cluster’s bulk motion. Neglecting these corrections can lead to artificial signatures, such as false rotation or contraction within the cluster. Correcting these projection-induced deviations is essential for accurately analyzing the intrinsic motion of stars.
To account for this, we corrected both the proper motions and radial velocities of stars by removing the projection-induced components, which depend on the cluster’s bulk motion, the star’s angular offset from the cluster center, and the parallax values of both the star and the cluster. The mathematical formalism for this correction is detailed in van Leeuwen (2009). The resulting corrected proper motions better reflect the intrinsic kinematics of the cluster.
4.2 Orbit Integration
Using the observed position and velocity data of the cluster, we computed its Galactic orbit employing the Astropy and Galpy libraries (Astropy Collaboration et al., 2013; Bovy, 2015; Astropy Collaboration et al., 2018, 2022). The orbit integration was carried out using the widely adopted MWPotential2014 model, which represents the Milky Way’s gravitational potential through a combination of an exponential disk, a spherical bulge, and a dark matter halo. We adopted the solar position $(-7999.986,0,15\leavevmode\nobreak\ \mathrm{pc})$ and velocity parameters $(10,235,7\leavevmode\nobreak\ \mathrm{km\,s^{-1}})$ provided in Bovy (2015).
Initial conditions for the orbit were defined using the cluster’s right ascension, declination, distance, proper motions, and radial velocity. The integration process involved numerically solving the equations of motion within the adopted Galactic potential. The resulting orbit was projected onto the tangent plane (the gray arrow in Figure 3) to visualize the cluster’s motion and identify the direction of its tidal tails. The leading tail points in the direction of orbital motion, while the trailing tail extends opposite to it.
@@ 4.3 Binary Fraction
Figure 3: First-order corrected proper motion vectors for the 5 open clusters.
@@ 5 Tidal Tails and their properties
This section presents results for the five clusters, focusing on tidal tails, internal dynamics, and orbits. We examined mass segregation, spatial and kinematic distributions, binary populations, and tail morphology.
To compare the tail and core populations, we applied the Kolmogorov–Smirnov (KS) test, reporting both the KS statistic and p-value: low p-values (<0.05<0.05) indicated significant differences, while high values suggested the samples are drawn from the same population. We also compared the observed properties with theoretical expectations and prior studies.
@@ 5.1 BH 164 (Alessi 7, MWSC 2255, vdBergh-Hagen 164)
Bhattacharya et al. (2022) identified 286 stars in BH 164, while our analysis found 396 stars within the cluster and its two tidal tails, which exhibit an S-shaped structure. The shape and orientation of the leading and trailing tails align with the theoretical N-body simulations (Figure 2), and the extent is consistent with Bhattacharya et al. (2022). Our study incorporated projection effects, identifying 289 cluster members and 107 tail stars, with the projected span of approximately 40 pc. In Figure (3), the proper motion vectors for cluster stars (blue) and tail stars (black and red) reveal a bifurcation pattern around the expected orbit, with opposing motions in the tails indicating outbound velocities.
Figure 7 (h) compares the luminosity functions (LFs) of the cluster (blue) and its tidal tails (red). The KS statistic of 0.10 and a p-value of 0.50 indicate no significant difference between the two distributions, implying that the tails contain a stellar population similar to the cluster. However, the cluster’s LF shows a steeper rise at the bright end, suggesting it retains a larger number of massive stars, while the tails are more populated by lower-mass stars. The binary fraction of the cluster was found to be $0.26\pm 0.04$, and the same average fraction of 0.26 is found for the tails. However, the binary fractions in the leading ($0.19\pm 0.07$) and trailing ($0.33\pm 0.09$) tails suggest an asymmetric ejection of binaries.
@@ 5.2 Alessi 2 (LeDrew 4, MWSC 419)
Bhattacharya et al. (2022) identified 201 stars in Alessi 2, whereas our analysis found 230 stars spanning the cluster and its two tidal tails, forming a subtle S -shaped structure along the expected orbit. The morphology and orientation are consistent with the theoretical predictions, but the N-body simulations suggested subtle S-shaped tails (Figure 2). After accounting for projection effects, we identified 130 core members and 100 tail stars, with a tip-to-tip extent of approximately 90 pc, significantly longer than the 47 pc reported by Bhattacharya et al. (2022).
Figure 3 shows that the tail stars exhibit no directional motion, with vectors pointing in various directions. The velocity vectors within the core appear to indicate motion predominantly in one direction, almost perpendicular to the expected orbit. However, upon closer examination, many of these vectors actually represent small motions in various directions, which are not readily visible in the projected plane plot of Figure 3.
Figure 8 (h) compares the luminosity functions (LFs) of the cluster and its tidal tails, with a KS statistic of 0.29 and p-value of 0.12. While the p-value suggests the distributions aren’t significantly different, the relatively high statistic points to a noticeable shift: the cluster retains higher-mass stars, while lower-mass stars dominate the tails. The binary fraction in the cluster was found to be $0.22\pm 0.05$, while the tidal tails exhibited a higher average fraction of $0.47$, with the leading and trailing tails having binary fractions of $0.54\pm 0.13$ and $0.40\pm 0.11$, respectively.
@@ 5.3 NGC 2281 (Collinder 116, MWSC 989, Melotte 51, OCL 446, Theia 635)
Tarricq et al. (2022) identified 443 stars in NGC 2281 with a tidal tail semi-major axis of $\approx$21 pc. Hunt & Reffert (2023, 2024) reported extended features as a byproduct of their analysis with projected extents based on our criteria of $\approx$34 pc and $\approx$19 pc for the leading and trailing tails, respectively, with a total tip-to-tip span of $\approx$80 pc.
The observed tail orientation is similar to the previous findings of Hunt & Reffert (2023, 2024), but we additionally identified a fragmented clump in the trailing tail. This clump at (-2, -4) deg in NGC 2281 appears to be separated from the rest of the tail candidates in both 2D and 3D space. Their location is different from what is expected from the simulations, and these stars are likely contaminants due to their comoving nature. Hence, we have manually removed these stars (crossed out in Figure 2). Our analysis after removing the clump detected 574 stars, including 438 cluster members and 136 tail stars. The leading and trailing tails span $\approx$39 pc and $\approx$25 pc, respectively, forming an S -shaped structure along the orbit, whose morphology and orientation aligned with the simulations (Figure 2). The total tip-to-tip extent came to $\approx$91 pc. Proper analysis of motion vectors (as shown in Figure 3) indicates that the leading and trailing tail vectors show similar motion with arrows pointing in opposite directions, leading to the elongation of tails. The motion of the cluster and the tail vectors are slightly angled relative to the direction of the orbital path.
@@ 5.4 NGC 2354 (Collinder 131, ESO 492-6, FSR 1291, MWSC 1148, OCL 639, Theia 5939)
Tarricq et al. (2022) detected 337 stars in NGC 2354 with tidal tails having a semi-major axis of $\approx$24 pc. The extended features reported by Hunt & Reffert (2023, 2024) based on our projected extent criteria span $\approx$22 pc (leading) and $\approx$9 pc (trailing), with a total tip-to-tip span of $\approx$61 pc.
Our analysis identified 386 stars, including 244 cluster members and 142 tail stars. The leading and trailing tails span $\approx$25 pc and $\approx$14 pc, respectively, with an overall tip-to-tip extent of $\approx$69 pc. The observed leading tail showed a local deviation, making it look S shaped. The orientation and morphology of the tidal tails in NGC 2354 (Figure 2) are consistent with the features reported by Hunt & Reffert (2023, 2024). The leading tail near (0, 1.5) degrees extends farther than predicted by the simulations, but in the absence of a clear discontinuity or any distinct break in the stellar distribution, we cannot rule out these stars as we did for NGC 2281. There is a possibility these tail candidates are false positives, and caution is advised before using the tails in NGC 2354. Proper motion vectors (Figure 3) reveal non-coherent and minimal motion in the cluster core as compared to the tails, which show large motion. The leading tail shows motion in the expected orbit’s direction, while the trailing tail moves in the reverse direction.
Figure 8 (h) shows the luminosity functions (LFs) of the cluster and its tidal tails, with a KS statistic of 0.21 and a p-value of 0.13. Although the p-value indicates that the difference is not statistically significant, the KS statistic suggests a moderate divergence. The cluster retains a higher number of massive stars, whereas the tails are dominated by lower-mass stars. The binary fraction in the cluster was found to be $0.34\pm 0.06$, compared to a lower average of $0.27$ in the tails. The leading and trailing tails show binary fractions of $0.23\pm 0.07$ and $0.31\pm 0.09$, respectively.

### v3

@@ 3.2 Color-magnitude diagram matching
3.3 Principal component analysis and DBSCAN clustering
@@
Figure 2: Spatial distribution of the identified core members (blue dots) for the five clusters with their detected tidal tails (red and coral dots), N-body simulated cluster (empty dots), Jacobi limit (dotted black circle), and expected orbit (gray arrow) pointing toward the direction of motion (refer to Subsect. 4.2). Crosses indicate manually removed stars for NGC 2281.
@@ 3.4 Tangential projection of the celestial coordinates
We employed a tangential projection of celestial coordinates onto a 2D Cartesian plane to analyze the spatial distribution of tidal tails in star clusters. This approach simplified visualization and enabled accurate spatial analyzes by mitigating the effects of spherical geometry.
For the trigonometric transformation, right ascension ($\alpha$) and declination ($\delta$) are first converted to radians, as is the reference point $(\alpha_{0},\delta_{0})$, corresponding to the mean position of the cluster core. The offsets in the tangent plane are computed as follows:
@@
The trigonometric expressions are dimensionless, but when interpreted under the small-angle approximation, they correspond to angular offsets expressed in radians. For visualization in Fig. 2, these offsets were explicitly converted to degrees. Although the projection provides a convenient 2D representation, it inevitably reduces information from the full 3D structure.
@@ 3.5 Beyond the Jacobi radius
The final criterion for checking our detection of the tidal tail was the spatial location of the stars relative to the best-fit $r_{J}$ values for each cluster as provided in Table 2 taken from Hunt & Reffert (2024), in which $r_{J}$ is defined as the distance from the cluster’s center to its $L_{1}$ Lagrange point, effectively marking the boundary beyond which stars are no longer bound to the cluster. Their full formulation accounts for the cluster mass as well as its orbital parameters around the host galaxy, including the circular frequency ($\Omega$) and the epicyclic frequency ($k$).
We note that $r_{J}$ is not used as a strict boundary in Fig. 2 for defining tidal tail stars. Due to projection effects in the sky plane, some tail stars may still appear within the projected $r_{J}$. Our use of this limit is only to show that the detected features extend beyond the theoretical boundary.
Table 2: Parameters of the five open clusters and their tidal tails. [^3] Col. 1: Cluster name; Cols. 2–4: Galactocentric positions of the identified members; Col. 5: Best-fitting $r_{J}$ from Hunt & Reffert (2024); Cols. 6–7: Approximate projected extent of the tidal tails beyond $r_{J}$ determined using the angular separation from the cluster center and the cluster distance, with the orbit direction aiding in identifying the tail components (Subsect. 4.2); Cols. 8-9: Number of stars in leading and trailing tails; Cols. 10-12: Binary fractions (for mass ratio >0.5>0.5) in the cluster, leading tail and trailing tail (Subsect. 4.4).
Name X Y Z $r_{J}$ $L_{\mathrm{lead}}$ $L_{\mathrm{trail}}$ $N_{\mathrm{lead}}$ $N_{\mathrm{trail}}$ $BF_{\mathrm{cluster}}$ $BF_{\mathrm{lead}}$ $BF_{\mathrm{trail}}$
@@
4 Tidal tail analysis
We analyzed the kinematic properties of each cluster and its tidal tails using proper motion vectors and orbit integration. Proper motions were used to characterize relative stellar motions and identify subtle patterns, while orbit integration traced past and future trajectories within the Milky Way’s potential. We also determined and analyzed the luminosity functions and the binary fractions for both clusters and their tails.
4.1 Correction of proper motion vectors for projection effects
Projection effects influence the apparent proper motions and RVs of individual stars within a star cluster due to the relative position and motion of the cluster with respect to the Sun (van Leeuwen, 2009; Jadhav et al., 2024). These effects arise because of the angular separation of stars from the cluster center and the cluster’s bulk motion. Neglecting these corrections can lead to artificial signatures, such as false rotation or contraction within the cluster. Correcting these projection-induced deviations is essential for accurately analyzing the intrinsic motion of stars.
To account for this, we corrected both the proper motions and RVs of stars by removing the projection-induced components, which depend on the cluster’s bulk motion, the star’s angular offset from the cluster center, and the parallax values of both the star and the cluster. The mathematical formalism for this correction is detailed in van Leeuwen (2009). The resulting corrected proper motions better reflect the intrinsic kinematics of the cluster.
4.2 Orbit integration
Using the observed position and velocity data of the cluster, we computed its Galactic orbit employing the Astropy and Galpy libraries (Astropy Collaboration et al., 2013; Bovy, 2015; Astropy Collaboration et al., 2018, 2022). The orbit integration was carried out using the widely adopted MWPotential2014 model, which represents the Milky Way’s gravitational potential through a combination of an exponential disk, a spherical bulge, and a dark matter halo. We adopted the solar position $(-7999.986,0,15\penalty 10000\ \mathrm{pc})$ and velocity parameters $(10,235,7\penalty 10000\ \mathrm{km\,s^{-1}})$ provided in Bovy (2015).
Initial conditions for the orbit were defined using the cluster’s ra, dec, distance, proper motions, and RV. The integration process involved numerically solving the equations of motion within the adopted Galactic potential. The resulting orbit was projected onto the tangent plane (the gray arrow in Fig. 2 and 3) to visualize the cluster’s motion and identify the direction of its tidal tails. The leading tail points in the direction of orbital motion, while the trailing tail extends opposite to it.
@@ 4.3 Luminosity function
The luminosity functions (LFs) for each cluster and its tidal tails were constructed to facilitate comparison, which can be seen in the subplot (h) of plots 7 to 11 in the appendix. Specifically, LFs were generated for two distinct regions: the identified core cluster and the detected tidal tails, including the leading tail and the trailing tail, each using a consistent bin size $M_{G}$ of 1 mag. The faintest one or two bins in both regions for every cluster contained fewer counts, probably due to observational incompleteness and stellar evaporation. These bins were excluded from the linear fit, which was performed using the following relation:
$\log_{10}N\propto a\times M_{G},$ (5)
@@ 4.4 Binary fraction
Figure 3: First-order corrected proper motion vectors for the five open clusters.
@@ 5 Tidal tails and their properties
This section presents results for the five clusters, focusing on tidal tails, internal dynamics, and orbits. We examined mass segregation, spatial and kinematic distributions, binary populations, and tail morphology. To compare the tail and core populations, we applied the Kolmogorov–Smirnov (KS) test, reporting both the KS statistic and p-value: low p-values (<0.05<0.05) indicated significant differences, while high values suggested the samples are drawn from the same population. We also compared the observed properties with theoretical expectations and prior studies.
@@ 5.1 BH 164 (Alessi 7, MWSC 2255, vdBergh-Hagen 164)
Bhattacharya et al. (2022) identified 286 stars in BH 164, while our analysis found 396 stars within the cluster and its two tidal tails, which exhibit an S-shaped structure. The shape and orientation of the leading and trailing tails align with the theoretical N-body simulations (Fig. 2), and the extent is consistent with Bhattacharya et al. (2022). Our study incorporated projection effects, identifying 289 cluster members and 107 tail stars, with the projected span of approximately 40 pc. In Fig. 3, the proper motion vectors for cluster stars (blue) and tail stars (black and red) reveal a bifurcation pattern around the expected orbit, with opposing motions in the tails indicating outbound velocities.
Figure 7 (h) compares the LFs of the cluster (blue) and its tidal tails (red). The KS statistic of 0.10 and a p-value of 0.50 indicate no significant difference between the two distributions, implying that the tails contain a stellar population similar to the cluster. However, the cluster’s LF shows a steeper rise at the bright end, suggesting it retains a larger number of massive stars, while the tails are more populated by lower-mass stars. The binary fraction of the cluster was found to be $0.26\pm 0.04$, and the same average fraction of 0.26 is found for the tails. However, the binary fractions in the leading ($0.19\pm 0.07$) and trailing ($0.33\pm 0.09$) tails suggest an asymmetric ejection of binaries.
@@ 5.2 Alessi 2 (LeDrew 4, MWSC 419)
Bhattacharya et al. (2022) identified 201 stars in Alessi 2, whereas our analysis found 230 stars spanning the cluster and its two tidal tails, forming a subtle S -shaped structure along the expected orbit. The morphology and orientation are consistent with the theoretical predictions, but the N-body simulations suggested subtle S-shaped tails (Fig. 2). After accounting for projection effects, we identified 130 core members and 100 tail stars, with a tip-to-tip extent of approximately 90 pc, significantly longer than the 47 pc reported by Bhattacharya et al. (2022). Figure 3 shows that the tail stars exhibit no directional motion, with vectors pointing in various directions. The velocity vectors within the core appear to indicate motion predominantly in one direction, almost perpendicular to the expected orbit. However, upon closer examination, many of these vectors actually represent small motions in various directions, which are not readily visible in the projected plane plot of Fig. 3.
Figure 8 (h) compares the LFs of the cluster and its tidal tails, with a KS statistic of 0.29 and p-value of 0.12. While the p-value suggests the distributions are not significantly different, the relatively high statistic points to a noticeable shift: the cluster retains higher-mass stars, while lower-mass stars dominate the tails. The binary fraction in the cluster was found to be $0.22\pm 0.05$, while the tidal tails exhibited a higher average fraction of $0.47$, with the leading and trailing tails having the binary fractions of $0.54\pm 0.13$ and $0.40\pm 0.11$, respectively.
@@ 5.3 NGC 2281 (Collinder 116, MWSC 989, Melotte 51, OCL 446, Theia 635)
Tarricq et al. (2022) identified 443 stars in NGC 2281 with a tidal tail semi-major axis of $\approx$21 pc. Hunt & Reffert (2023, 2024) reported extended features as a byproduct of their analysis with projected extents based on our criteria of $\approx$34 pc and $\approx$19 pc for the leading and trailing tails, respectively, with a total tip-to-tip span of $\approx$80 pc. The observed tail orientation is similar to the previous findings of Hunt & Reffert (2023, 2024), but we additionally identified a fragmented clump in the trailing tail. This clump at (-2, -4) deg in NGC 2281 appeared to be separated from the rest of the tail candidates in both 2D and 3D space. The location is also different from what is expected from the simulations, and these stars are likely contaminants due to their comoving nature. Hence, we have manually removed these stars (crossed out in Fig. 2). Our analysis after removing the clump detected 574 stars, including 438 cluster members and 136 tail stars. The leading and trailing tails span $\approx$39 pc and $\approx$25 pc, respectively, forming an S -shaped structure along the orbit, whose morphology and orientation aligned with the simulations (Fig. 2). The total tip-to-tip extent came to $\approx$91 pc. Proper analysis of motion vectors (as shown in Fig. 3) indicates that the leading and trailing tail vectors show similar motion with arrows pointing in opposite directions, leading to the elongation of tails. The motion of the cluster and the tail vectors are slightly angled relative to the direction of the orbital path.
@@ 5.4 NGC 2354 (Collinder 131, ESO 492-6, FSR 1291, MWSC 1148, OCL 639, Theia 5939)
Tarricq et al. (2022) detected 337 stars in NGC 2354 with tidal tails having a semi-major axis of $\approx$24 pc. The extended features reported by Hunt & Reffert (2023, 2024) based on our projected extent criteria span $\approx$22 pc (leading) and $\approx$9 pc (trailing), with a total tip-to-tip span of $\approx$61 pc. Our analysis identified 386 stars, including 244 cluster members and 142 tail stars. The leading and trailing tails span $\approx$25 pc and $\approx$14 pc, respectively, with an overall tip-to-tip extent of $\approx$69 pc. The observed leading tail showed a local deviation, making it look S shaped. The orientation and morphology of the tidal tails in NGC 2354 (Fig. 2) are consistent with the features reported by Hunt & Reffert (2023, 2024). The leading tail near (0, 1.5) deg extended farther than predicted by the simulations, but in the absence of a clear discontinuity or any distinct break in the stellar distribution, we could not rule out these stars as we did for NGC 2281. There is a possibility these tail candidates are false positives, and caution is advised before using the tails in NGC 2354. Proper motion vectors (Fig. 3) reveal noncoherent and minimal motion in the cluster core as compared to the tails, which show large motion. The leading tail shows motion in the expected orbit’s direction, while the trailing tail moves in the reverse direction.
Figure 10 (h) shows the LFs of the cluster and its tidal tails, with a KS statistic of 0.21 and a p-value of 0.13. Although the p-value indicates that the difference is not statistically significant, the KS statistic suggests a moderate divergence. The cluster retains a higher number of massive stars, whereas the tails are dominated by lower-mass stars. The binary fraction in the cluster was found to be $0.34\pm 0.06$, compared to a lower average of $0.27$ in the tails. The leading and trailing tails show the binary fractions of $0.23\pm 0.07$ and $0.31\pm 0.09$, respectively.